OpenStudy (anonymous):

why does the square root of 2 over 2 + the square root of 2 over 2 =the square root of 2?

OpenStudy (anonymous):

a/2 + a/2 = a a = square root of 2

OpenStudy (anonymous):

\[\sqrt{2}/2 + \sqrt{2}/2 = \sqrt{2} (1/2+1/2)= \sqrt{2} (1) = \sqrt{2}\]

OpenStudy (anonymous):

Think of getting the sqrt of 2 out, as a common factor. Then add whats inside the bracket and then what do you get? :)

OpenStudy (anonymous):

\[\sqrt{2}/2 = \sqrt{2} * 1/2 \]